OpenStudy (anonymous):

What is the inverse of f(x) = 2x − 8?

OpenStudy (anonymous):

Inverse are.... you replace the x's with y's and the y's with x's y = 2x - 8 is your equation x = 2y - 8 is the inverse

OpenStudy (anonymous):

thats not complete. you switch the X and Y then solve for Y. then you get ure inverse

OpenStudy (anonymous):

I dunno why people always make inverses so odd.. \[f(x)=2x-8 \text{ , } f(f^{-1}(x)) = x \] \[\implies x = 2f^{−1}(x) - 8\] \[\implies 2f^{−1}(x) = x + 8\] \[\implies f^{−1}(x) = \frac{1}{2}x + 4\]